OpenStudy (anonymous):

find the vector and the cartesian equation of the line passing through (1,2,3) and parallel to the planes x-y+2z=5 and 3x+y+z=6

OpenStudy (anonymous):

@Sauravshakya

OpenStudy (experimentx):

let l,m,n be the direction cosines of your line, the line is parallel to the planes which means the line is perpendicular to the dc of the normal of the line.

OpenStudy (anonymous):

would be plse explain more or do the sum fro me..because i have some problem in these sums.

OpenStudy (experimentx):

when two vectors are perpendicular, their dot product is zero. the set of direction cosines of lines <l,m,n> and the set of direction cosines of normal of plane <a,b,c> are perpendicular.

OpenStudy (experimentx):

find the direction cosines of the normal of the plane/

OpenStudy (anonymous):

Plane parallel to x-y+2z=5 and that passes through (1,2,3) is x-y+2z=5 ....i Plane parallel to3x+y+z=6 and that passes through (1,2,3) is 3x+y+z=8 ...ii

OpenStudy (anonymous):

Now, u need to find equation of line i.e intersection of plane i and ii

OpenStudy (experimentx):

works perfectly that way too ... since a line is a intersection of 2 planes in 3d

OpenStudy (anonymous):

@sauravshakya , plse continue

OpenStudy (anonymous):

x-y+2z=5.....i 3x+y+z=8 ....ii Let x=0 then -y+2z=5 ..iii y+z=8...iv This gives z=13/3 and y=11/3 So, one of the coordinate of point that lies in the line is (0,11/3 ,13/3)

OpenStudy (anonymous):

Similarly, Let z=0 and u will get x=13/4 and y=-7/4 So, another coordinate of the point is (13/4 , -7/4,0)

OpenStudy (anonymous):

Now, can u do the rest?

OpenStudy (anonymous):

srry can u plse do further also..

OpenStudy (experimentx):

use the definition of the equation of line http://answers.yahoo.com/question/index?qid=20080830195656AA3aEBr

OpenStudy (anonymous):

i m not getting it..

OpenStudy (anonymous):

@Sauravshakya

OpenStudy (anonymous):

can anyone help me

OpenStudy (anonymous):

@AravindG

OpenStudy (experimentx):

 this is the equation of your line. x0,y0,z0 and x1,y1,z1 are your two points.